A lump of gold has a density of 6gcm3 of volume of 20cm3. Calculate the mass of the gold
1
Expert's answer
2018-06-20T09:58:09-0400
By definition density is given by
ρ=m/V,
where m – is the mass of the body, and V – occupied volume.
Therefore
m=ρ∙V
Since it’s given that ρ=6 g/(cm^3 ) and V=20cm^3, we obtain
m=6 g/(cm^3 )∙20cm^3=120g